When I was at your age now, you were only 5 years old. When you become as old as I am now, I will be 71 years old." How old are A and B now, respectively?
Found 2 solutions by ikleyn, ankor@dixie-net.com:
Answer by ikleyn(52839)

.
"When I was at your age now, you were only 5 years old. When you become as old as I am now, I will be 71 years old." 
How old are A and B now, respectively?
---------------------------------------------------------------

Let me re-formulate to get the solution more easily.

===============================================================
When A was at the age of B now, B was only 5 years old. When B become as old as A is now, A will be 71 years old. 
How old are A and B now, respectively?
===============================================================

Let x be the preset age of A and y be the present age of B.

When A was at the age of B now? - It was x-y years ago. 
What was the age of B at that time, x-y years ago? The age of B was y - (x-y) = 2y-x. 

It gives you the first equation

2y - x = 5.

When B become as old as A is now? - It will happen after x-y years.
What will be the age of A at that time? - The age of A at that time will be x + (x-y) = 2x - y. 

It gives you the second equation

2x - y = 71.

Now solve this system

2y - x = 5,
2x - y = 71.

Do it yourself please. It is easy.

Answer. A is 49 years now. The present age of B is 27 years.


Answer by ankor@dixie-net.com(22740)

When I was at your age now, you were only 5 years old.
When you become as old as I am now, I will be 71 years old."
How old are A and B now, respectively?
:
let d = the difference in their ages
d = A - B
"When I was at your age now, you were only 5 years old.
B - d = 5
B = d + 5
"When you become as old as I am now, I will be 71 years old."
A + d = 71
A = 71 - d
:
d = A - B
Replace A with (71-d); replace with (d+5)
d = (71-d) - (d+5)
d = 71 - d - d - 5
d = 66 - 2d
d + 2d = 66
3d = 66
d = 22 yrs is the difference in their ages
Find A
A = 71 - d
A = 71 - 22
A = 49 yrs is A's age
Find B
B = d + 5
B = 22 + 5
B = 27 yrs is B's age